Can anyone help me with ANY of these? I don't understand how the roots come into play. The title of this section is Sum and Product of Roots.
Question#1
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
a = __ b = __ c = __

Question #2

x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(-b/a)= -__ / __

Question #3
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 + root2 = 5 + -11 = ___

Question #4

x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The sum of the roots is the same as (Hint: use letters)
= -__ /__

Question #5
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(c/a) = __ /__
Question #6
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 x root2 = 5 x (-11) = ___
Question #7
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The product of the roots is the same as (Hint: use letters)
= - ___ /___
Question #8
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
Identify a, b, and c. (Hint: you may have to arrange the equation in the form ax2+ bx + c = 0 .)
a =___ b = ___ c = ___
Question #9
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(-b/a)= -___ /___
Question #10
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 + root2 =(1/2)+(-1/4)= ___
Question #11

8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The sum of the roots is the same as (Hint: use letters)
= ____

Question #12
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(c/a)= ____ /___
Question #13
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 x root2 = (1/2)(-1/4) = ___

Question #14
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The product of the roots is the same as (Hint: use letters)
= ___ /___

Answer by stanbon(75887)

Question#1
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
a = _1_ b = _6_ c = _-55_
Question #2
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(-b/a)= -_6_ / _1_ = -6
Question #3
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 + root2 = 5 + -11 = _-6__
Question #4
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The sum of the roots is the same as (Hint: use letters)
= -_ba_ /__
Question #5
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(c/a) = _-55_ /_1_ = -55
Question #6
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 x root2 = 5 x (-11) = _-55__
Question #7
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The product of the roots is the same as (Hint: use letters)
= - __c_ /_a__
Question #8
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
Identify a, b, and c. (Hint: you may have to arrange the equation in the form ax2+ bx + c = 0 .)
a =__8_ b = _-2__ c = _-1__
Question #9
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(-b/a)= -__-2_ /_8__ = 1/4
Question #10
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 + root2 =(1/2)+(-1/4)= __1/4--_

Question #11
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The sum of the roots is the same as (Hint: use letters)
= _-b/a___
Question #12
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(c/a)= _-1___ /_8__ = -1/8
Question #13
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 x root2 = (1/2)(-1/4) = _-1/8__
Question #14
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The product of the roots is the same as (Hint: use letters)
= _c__ /_a__
=========================
Cheers,
Stan H.


--------------------------------------------------------------------------------

Textbook information not provided.What's this?
Quadratic_Equations/202089 (2009-07-05 18:11:37): Can anyone help me with ANY of these? I don't understand how the roots come into play. The title of this section is Sum and Product of Roots.
Question#1
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
a = __ b = __ c = __
Question #2
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(-b/a)= -__ / __
Question #3
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 + root2 = 5 + -11 = ___
Question #4
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The sum of the roots is the same as (Hint: use letters)
= -__ /__
Question #5
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
(c/a) = __ /__
Question #6
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
root1 x root2 = 5 x (-11) = ___
Question #7
x2 + 6x -55 = 0 and root1 = 5, root2 = -11
The product of the roots is the same as (Hint: use letters)
= - ___ /___
Question #8
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
Identify a, b, and c. (Hint: you may have to arrange the equation in the form ax2+ bx + c = 0 .)
a =___ b = ___ c = ___
Question #9
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(-b/a)= -___ /___
Question #10
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 + root2 =(1/2)+(-1/4)= ___
Question #11
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The sum of the roots is the same as (Hint: use letters)
= ____
Question #12
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
(c/a)= ____ /___
Question #13
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
root1 x root2 = (1/2)(-1/4) = ___
Question #14
8x2 - 2x = 1 and root1= 1/2 root2= -1/4
The product of the roots is the same as (Hint: use letters)
= ___ /___